# HRNet: Differentially Private Hierarchical and Multi-Resolution Network for Human Mobility Data Synthesization

### v1

@@ Abstract.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
@@
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/tkgsn/priv_traj_gen.
@@ 1. Introduction
Differential Privacy (DP) (Dwork, 2006) has emerged as the leading standard for maintaining data privacy. DP offers a strong, mathematically grounded privacy guarantee without relying on restrictive assumptions about potential adversaries. The core principle of DP is to ensure that output from the data analysis does not substantially differ, regardless of whether any specific individual’s data is included or excluded from the dataset. The broad applicability and growing adoption of DP in the real world (Erlingsson et al., 2014; Team, 2017; Ding et al., 2017; Johnson et al., 2018), as well as its endorsement by the US Census Bureau (Bureau, 2020), signify its importance and effectiveness.
@@
In this paper, we study the problem of DP data synthesis (Zhang et al., 2017; Bowen and Liu, 2021; Li et al., 2021) for human mobility data. DP data synthesis generates synthetic data that possesses statistical properties similar to the real data while ensuring DP. According to the post-processing theorem of DP, unlimited analysis can be performed on this synthetic data without introducing further privacy concerns. Our goal is to synthesize human mobility data, defined as a sequence of locations, to closely resemble the provided human mobility data, which can be then used for a variety of downstream tasks in previously mentioned applications.
@@
Existing DP data synthesis methods for human mobility primarily rely on DP-aware data structures, which can be categorized into tree-based (Chen et al., 2012; He et al., 2015; Cai et al., 2021), Markov-based (Wang et al., 2023; Gursoy et al., 2018), and clustering approaches (Liu et al., 2021; Hua et al., 2015; Li et al., 2017). These methods generally employ generalization to reduce dimensionality and sensitivity, thereby balancing utility and privacy, but often at the cost of information loss. For instance, clustering approaches only generate transitions between $k$ centroids by location generalization using clustering, leading to a loss of precise transition information within the centroids. Moreover, the state-of-the-art Markov-based method (Wang et al., 2023) does not account for correlations beyond two-step intervals, and many methods overlook auxiliary information such as time, which is crucial in human movement. Tree-based methods aim to minimize such loss by constructing a data-dependent tree with a portion of the privacy budget. However, such consumption leads to a lack of privacy budget during the data generation phase, which results in diminished utility.
@@
MTNet, proposed by Wang et al. (Wang et al., 2022), is a deep learning approach designed for generating short trips on road networks, such as taxi trips, while ensuring DP. By utilizing meta-information from road networks, which is not sensitive, MTNet adeptly manages the utility-privacy trade-off. However, this approach imposes the inherent limitation of restricting mobility to road networks. Human mobility, in reality, is not confined to such networks; it encompasses various modes including subway travel and movement through non-road spaces like parks. Additionally, even when focused on road networks, human mobility often evolves into lengthy sequences of road segments, a complexity that poses significant challenges under DP constraints. Consequently, adapting MTNet for a broader spectrum of human mobility appears to be an unpromising approach.
@@
To the best of our knowledge, this is the first paper to adopt a deep learning approach with DP for generating human mobility which is not constrained to road networks. Recent advancements have shown that deep learning approaches handling human mobility data outperform traditional models (Wang et al., 2020; Luca et al., 2023; Zhang et al., 2023). There are several generative models such as adversarial network (GAN) based or diffusion models for generating realistic synthetic trajectories. However, these models do not ensure formal privacy guarantee. Theoretically, it is possible to ensure DP for any deep learning models by replacing the optimization method with differential private stochastic gradient descent (DP-SGD) (Abadi et al., 2016). Unfortunately, simply adopting DP-SGD on these generative models results in either poor utility or infeasibly high privacy cost due to the large space domain and complexity of the models. More concretely, we elaborate the two primary challenges as follows.
@@
    The number of model parameters: It has been demonstrated (Bassily et al., 2014, 2020) that DP-SGD inevitably increases the lower bound of empirical risk as the number of parameters increases. In most models about human mobility (e.g., an RNN model (Yang et al., 2020) and an attention model (Yan et al., 2023)), the number of model parameters depends on the size of the space domain. If we encode the space or Points of Interests (POIs) using an encoding scheme such as embedding matrices or linear embeddings, the number of model parameters is $O(n_{\rm POI})$ with respect to $n_{\rm POI}$ where $n_{\rm POI}$ denotes number of POIs, which can be significantly large.
@@
    Learning difficulty: the difficulty of learning human mobility escalates with an increase in $n_{\rm POI}$ because of the increasing number of labels for training. Therefore, a larger $n_{\rm POI}$ necessitates larger models and longer training epochs to discern subtle POI differences. This, in turn, leads to decreased utility due to the large number of parameters and higher consumption of privacy budgets due to the composition theorem (Kairouz et al., 2015).
@@ 2.1. Differential Privacy
DP-SGD: Differentially Private Stochastic Gradient Descent (DP-SGD), as developed by Abadi et al. (Abadi et al., 2016), adapts the conventional SGD optimization process for deep learning models to satisfy DP. In traditional SGD, parameters of a model $\theta$ are iteratively updated to minimize empirical risk, with gradient computations performed using sampled data (minibatches). DP-SGD modifies this process to ensure DP by introducing gradient clipping and noise addition steps. Specifically, the $l_{2}$ norm of the gradient is first clipped to limit sensitivity, followed by the addition of the Gaussian noise to the averaged clipped gradients. The model parameters are then updated using these randomized gradients, similar to conventional SGD. This process continues until convergence is achieved or the allocated privacy budget is depleted. From a DP perspective, this process is interpreted as the sequential composition of the Gaussian mechanism with subsampling (i.e., minibatch). Therefore, by employing composition techniques, the final published model satisfies DP. We adopt numerical composition, which numerically computes the upper bound of the privacy loss parameter $\varepsilon$ using privacy loss distributions (Sommer et al., 2018). Several studies enhance its computation using techniques such as the Fast Fourier Transform (Koskela et al., 2020) and the ”connect-the-dots” algorithm (Dockhorn et al., 2022). For implementation, we utilize Google’s differential-privacy library  [^1] https://github.com/google/differential-privacy.
@@ 2.2.1. Trajectory:
As a fundamental step, we define Points of Interests (POIs) [^2] Here, the term ’POI’ is not used in its original sense of indicating semantic locations. However, our method can be applied in its original context as described in Section 3.2.2. as uniformly distributed grid cells on a map, with each grid cell assigned a unique integer ID. For instance, if a target map is divided into $w\times w$ grids, the domain of POIs is defined as $L:=[w^{2}]$, where $[n]$ denotes the set $(1,2,\dots,n)$. That is, $n_{\rm POI}=|L|$. As illustrated in Figure 1, different levels of discretization are possible, such as $1\times 1$, $2\times 2$, and $4\times 4$. A geographical location, defined by its latitude and longitude, is transformed into an integer representing the grid cell that encompasses the location. The mapping is performed such that the grid located at $(\lfloor l/w\rfloor,l\%w)$ is assigned the integer $l\in L$. We then represent a human mobility sequence as a sequence of stay points, denoted as $\mathbf{v}=[v_{1},v_{2},\dots],$ where $i\in\mathbb{N},v_{i}=(l_{i},t_{i})$, $l_{i}\in L$. Note that this representation aligns POIs that change the previous POI so that $l_{i}\neq l_{i+1}$. Here, $t_{i}$ is the index of the time slot that includes the corresponding time. Time slots are made by discretizing entire time by $n_{\rm time}$, where $n_{\rm time}$ is the number of time slots. For example, $v_{2}$ indicates the tuple of the second POI visited and the time at which this POI was reached. From this point forward, the term ’trajectory’ will refer to this discretized representation of human mobility.
@@ 2.2.3. Evaluation:
Direct evaluation of the performance of $G_{\theta}$ is intractable due to the inherent sparsity of trajectory data. To overcome this, we adopt an indirect evaluation approach, focusing on various statistical properties of trajectories. These statistical properties are computed from the dataset generated by $G_{\theta}$ and are then compared with empirical statistics derived from the provided trajectory dataset $D$. Key aspects of trajectory data that we examine include waypoints, routes, destinations, transitions, and travel distances. By analyzing these specific statistics, we can gain insights into the accuracy and fidelity of the trajectories generated by $G_{\theta}$. The formal definitions and methodologies for computing these statistics are comprehensively detailed in Section 5.1. This approach enables a thorough and nuanced evaluation of the generator’s performance in replicating realistic human mobility patterns.
@@ 2.3.1. The baseline generator
Recurrent Neural Network: For encoding trajectory prefixes $(v_{1},\dots,v_{i-1})$, we employ Gated Recurrent Units (GRUs) (Cho et al., 2014). This is because we have empirically found that GRU is slightly more effective than alternative methods such as LSTM and attention mechanisms in the context of DP-SGD. We attribute this to the GRU’s simpler architecture and reduced parameter. The GRU cell updates its state based on the previous state $h_{i-1}$ and the encoded $v_{i-1}$:
$h_{i}=f_{\rm GRU}(h_{i-1},{\rm encode}(v_{i-1})).$
This process is applied recursively to encode the entire trajectory prefix.
Scoring component: The scoring component is responsible for converting the encoded prefix into the probability distribution over $[n_{\rm POI}]\times[n_{\rm time}]$. This conversion is accomplished using feed-forward neural networks $g_{\rm POI}:\mathbb{R}^{n_{\rm hidden}}\to\mathbb{R}^{|L|}$ and $g_{\rm time}:\mathbb{R}^{n_{\rm hidden}}\to\mathbb{R}^{n_{\rm time}}$ where $n_{\rm hidden}$ is the dimensionality of the hidden state $h_{i}$, followed by a softmax function to ensure proper probabilistic normalization. The probability distribution over spatial location $l\in L$ given the hidden state $h_{i}$ is computed as $\Pr(l=l_{i}|h_{i})={\rm softmax}(g_{\rm POI}(h_{i}))[l_{i}].$ Similarly, the probability distribution over time $t\in[n_{\rm time}]$ given $h_{i}$ is determined by $\Pr(t=t_{i}|h_{i})={\rm softmax}(g_{\rm time}(h_{i}))[t_{i}].$ Ultimately, the conditional probability $\Pr(v_{i}|v_{1},\dots,v_{i-1})$ given the previous observations is the product of these two distributions, representing a joint probability over POI and time: $\Pr(v_{i}|v_{1},\dots,v_{i-1})=\Pr(l=l_{i}|h_{i})\Pr(t=t_{i}|h_{i}).$
@@ 2.3.2. The objective function:
To optimize the baseline generator, a given trajectory sequence $\mathbf{v}=(v_{1},\dots,v_{i})$, we employ cross entropy loss as our objective function. The loss is calculated as follows:
@@ 2.4. The two bottlenecks
Under the constraints of DP, it is known that the lower bound of empirical risk of the private model polynomially scales (Bassily et al., 2014, 2020) as the number of parameters of the model increases. In the baseline model, the size of the parameter vector $\theta$ expands linearly with the increasing number of POIs ($n_{\rm POI}$) due to the location embedding matrix and the scoring component. Consequently, the empirical risk would polynomially escalate as $n_{\rm POI}$ increases, which leads to worse utility of the generated human mobility.
@@ 2.4.2. The number of labels
With an increase in the number of class labels (i.e., $n_{\rm POI}$), the model faces challenges in discerning more subtle feature distinctions between classes (see Section 2.3.2). This often necessitates the use of larger model architectures or extended training epochs. However, such strategies are impractical under DP-SGD constraints. This is because an increase in the number of parameters in the model leads to an increase in empirical risk, as mentioned above (Bassily et al., 2014, 2020). Additionally, increasing the number of training epochs results in greater privacy loss due to the composition theorem of DP (Kairouz et al., 2015). Therefore, learning difficulty caused by the large number of labels is a bottleneck of DP-SGD.
@@ 2.5. Transposed Convolution
Consider the matrix $M$ with shape $(w,w,n_{\rm dim})$, where $n_{\rm dim}$ denotes the number of channels, and each channel comprises a $w\times w$ square matrix. Similarly, let the deconvolutional kernel be defined as a matrix of shape $(2,2,n_{\rm dim})$, containing $n_{\rm dim}$ kernels, and each of these kernels is $2\times 2$ matrix denoted by ${\rm kernel}_{k}$. The 2D quad deconvolutional network expands the matrix $M$ to the matrix $M^{\prime}$ with dimensions $(2w,2w,n_{\rm dim})$ using a deconvolutional kernel. See Figure 1 for the running example. In the second layer, the kernel slides over the input of $(2\times 2)$ matrix to perform the deconvolution operation for the four grids, expanding a $2\times 2$ matrix into a $4\times 4$ matrix

### v2

@@ Abstract.
* Work partially done while visiting Emory University
@@
Proceedings of the VLDB Endowment, Vol. 17, No. 11 ISSN 2150-8097.
@@
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/Emory-AIMS/priv_traj_gen.
@@ 1. Introduction
Differential Privacy (DP) (Dwork, 2006) has emerged as the leading standard for maintaining data privacy. DP offers a strong, mathematically grounded privacy guarantee without relying on restrictive assumptions about potential adversaries. The core principle of DP is to ensure that output from the data analysis does not substantially differ (bounded by a privacy parameter or privacy budget), regardless of whether any specific individual’s data is included or excluded from the dataset. The broad applicability and growing adoption of DP in the real world (Erlingsson et al., 2014; Team, 2017; Ding et al., 2017; Johnson et al., 2018), as well as its endorsement by the US Census Bureau (Bureau, 2020), signify its importance and effectiveness.
@@
In this paper, we study the problem of DP data synthesis (Zhang et al., 2017; Bowen and Liu, 2021; Li et al., 2021) for human mobility data. DP data synthesis generates synthetic data that possesses statistical properties similar to the real data while ensuring DP. According to the post-processing theorem of DP, unlimited analysis can be performed on this synthetic data without introducing further privacy concerns. Our goal is to synthesize human mobility data, defined as a sequence of locations, to closely resemble the real human mobility data, which can be then used for a variety of downstream tasks in previously mentioned applications.
@@
Existing DP data synthesis methods for human mobility primarily rely on DP-aware data structures, which can be categorized into tree-based (Chen et al., 2012; He et al., 2015; Cai et al., 2021), Markov-based (Wang et al., 2023; Gursoy et al., 2018), and clustering approaches (Liu et al., 2021; Hua et al., 2015; Li et al., 2017). These methods generally employ generalization to reduce dimensionality and sensitivity, thereby balancing utility and privacy, but often at the cost of information loss. For instance, clustering approaches only generate transitions between $k$ centroids by location generalization using clustering, leading to a loss of precise transition information within the clusters. Moreover, the state-of-the-art Markov-based method (Wang et al., 2023) does not account for correlations beyond two-step intervals, and many methods overlook auxiliary information such as time, which is crucial in human movement. Tree-based methods aim to minimize such loss by constructing a data-dependent tree with a portion of the privacy budget. However, such consumption leads to a lack of privacy budget during the data generation phase, which results in diminished utility.
@@
MTNet (Wang et al., 2022) is a deep learning approach designed for generating short trips on road networks, such as taxi trips, while ensuring DP. By utilizing meta-information from road networks, which is not sensitive, MTNet adeptly manages the utility-privacy trade-off. However, this approach imposes the inherent limitation of restricting mobility to road networks. Human mobility, in reality, is not confined to such networks; it encompasses various modes including subway travel and movement through non-road spaces like parks. Additionally, even when focused on road networks, human mobility often evolves into lengthy sequences of road segments, a complexity that poses significant challenges under DP constraints. Consequently, adapting MTNet for a broader spectrum of human mobility appears to be an unpromising approach.
@@
To the best of our knowledge, this is the first paper to adopt a deep learning approach with DP for generating human mobility which is not constrained to road networks. Recent advancements have shown that deep learning approaches handling human mobility data outperform traditional models (Wang et al., 2020; Luca et al., 2023). There are several generative models such as adversarial network (GAN) based models for generating realistic synthetic trajectories (Zhang et al., 2023; Feng et al., 2020). However, these models do not ensure formal privacy guarantee. Theoretically, it is possible to ensure DP for any deep learning models by replacing the optimization method with differentially private stochastic gradient descent (DP-SGD) (Abadi et al., 2016). Unfortunately, simply adopting DP-SGD on these generative models results in either poor utility or infeasibly high privacy cost due to the large space domain and complexity of the models. More concretely, we elaborate the two primary challenges as follows.
@@
    The number of model parameters: It has been demonstrated (Bassily et al., 2014, 2020) that DP-SGD inevitably increases the lower bound of empirical risk as the number of parameters increases. In most models about human mobility (e.g., RNN model (Yang et al., 2020) and attention model (Yan et al., 2023)), the number of model parameters depends on the size of the space domain. If we encode the space or Points of Interests (POIs) using an encoding scheme such as embedding matrices or linear embeddings, the number of model parameters is $O(n_{\rm POI})$ where $n_{\rm POI}$ denotes number of POIs, which can be significantly large.
@@
    Learning difficulty: the difficulty of learning human mobility escalates with increasing $n_{\rm POI}$ due to the increasing number of labels for training. Therefore, a larger $n_{\rm POI}$ necessitates larger models and longer training epochs to discern subtle POI differences. This leads to decreased utility due to the large number of parameters and higher consumption of privacy budgets due to the composition theorem (Kairouz et al., 2015).
@@ 2.1. Differential Privacy
DP-SGD: Differentially Private Stochastic Gradient Descent (DP-SGD), as developed by Abadi et al. (Abadi et al., 2016), adapts the conventional SGD optimization process for deep learning models to satisfy DP. In traditional SGD, parameters of a model $\theta$ are iteratively updated to minimize empirical risk, with gradient computations performed using sampled data (minibatches). DP-SGD modifies this process to ensure DP by introducing gradient clipping and noise addition steps. Specifically, the $l_{2}$ norm of the gradient is first clipped to limit sensitivity, followed by the addition of the Gaussian noise to the averaged clipped gradients. The model parameters are then updated using these randomized gradients, similar to conventional SGD. This process continues until convergence is achieved or the allocated privacy budget is depleted. From a DP perspective, this process is interpreted as the sequential composition of the Gaussian mechanism with subsampling (i.e., minibatch). Therefore, by employing composition techniques, the final published model satisfies DP. We adopt numerical composition, which numerically computes the upper bound of the privacy loss parameter $\varepsilon$ using privacy loss distributions (Sommer et al., 2018). Several studies enhance its computation using techniques such as the Fast Fourier Transform (Koskela et al., 2020) and the ”connect-the-dots” algorithm (Dockhorn et al., 2022). For implementation, we utilize Google’s differential-privacy library [^1] https://github.com/google/differential-privacy (Accessed: July 14, 2024).
@@ 2.2.1. Trajectory:
As a fundamental step, we define Points of Interests (POIs) [^2] Here, the term ‘POI’ is not used in its original sense of indicating semantic locations. However, our method can be applied in its original context as described in Section 3.2.2. as uniformly distributed grid cells on a map, with each grid cell assigned a unique integer ID. For instance, if a target map is divided into a $w\times w$ grid, the domain of POIs is defined as $L:=[w^{2}]$, where $[n]$ denotes the set $(1,2,\dots,n)$. That is, $n_{\rm POI}=|L|$. As illustrated in Figure 1, different levels of discretization are possible, such as $1\times 1$, $2\times 2$, and $4\times 4$. A geographical location, defined by its latitude and longitude, is transformed into an integer representing the grid cell that encompasses the location. The mapping is performed such that the grid cell located at $(\lfloor l/w\rfloor,l\%w)$ is assigned the integer $l\in L$. We then represent a human mobility sequence as a sequence of stay points, denoted as $\mathbf{v}=[v_{1},v_{2},\dots],$ where $i\in\mathbb{N},v_{i}=(l_{i},t_{i})$, $l_{i}\in L$. Note that this representation aligns POIs that change from the previous POI so that $l_{i}\neq l_{i+1}$. Here, $t_{i}$ is the index of the time slot that includes the corresponding time. Time slots are made by discretizing entire time by $n_{\rm time}$, where $n_{\rm time}$ is the number of time slots. For example, $v_{2}$ indicates the tuple of the second POI visited and the time at which this POI was reached. From this point forward, the term ‘trajectory’ will refer to this discretized representation of human mobility.
@@ 2.2.3. Evaluation:
Direct evaluation of the performance of $G_{\theta}$ is intractable due to the inherent sparsity of trajectory data. To overcome this, we adopt an indirect evaluation approach, focusing on various statistical properties of trajectories. These statistical properties are computed from the dataset generated by $G_{\theta}$ and are then compared with empirical statistics derived from the provided trajectory dataset $D$. Key aspects of trajectory data that we examine include waypoints, routes, destinations, transitions, and travel distances. By analyzing these specific statistics, we can gain insights into the accuracy and fidelity of the trajectories generated by $G_{\theta}$. The formal definitions and methodologies for computing these statistics are detailed in Section 5.1. This approach enables a thorough and nuanced evaluation of the generator’s performance in replicating realistic human mobility patterns.
@@ 2.3.1. The baseline generator
Recurrent Neural Network: For encoding trajectory prefixes $(v_{1},\dots,v_{i-1})$, we employ Gated Recurrent Units (GRUs) (Cho et al., 2014). This is because we have empirically found that GRU is slightly more effective than alternative methods such as LSTM (Hochreiter and Schmidhuber, 1997) and attention mechanisms (Vaswani et al., 2017) in the context of DP-SGD. We attribute this to the GRU’s simpler architecture and reduced parameters. The GRU cell updates its state based on the previous state $h_{i-1}$ and the encoded $v_{i-1}$: $h_{i}=f_{\rm GRU}(h_{i-1},{\rm encode}(v_{i-1})).$ This process is applied recursively to encode the entire trajectory prefix.
Scoring component: The scoring component is responsible for converting the encoded prefix into the probability distribution over $[n_{\rm POI}]\times[n_{\rm time}]$ for next predicted location. This conversion is accomplished using feed-forward neural networks $g_{\rm POI}:\mathbb{R}^{n_{\rm hidden}}\to\mathbb{R}^{|L|}$ and $g_{\rm time}:\mathbb{R}^{n_{\rm hidden}}\to\mathbb{R}^{n_{\rm time}}$ where $n_{\rm hidden}$ is the dimensionality of the hidden state $h_{i}$, followed by a softmax function to ensure proper probabilistic normalization. The probability distribution over spatial location $l\in L$ given the hidden state $h_{i}$ is computed as $\Pr(l=l_{i}|h_{i})={\rm softmax}(g_{\rm POI}(h_{i}))[l_{i}].$ Similarly, the probability distribution over time $t\in[n_{\rm time}]$ given $h_{i}$ is determined by $\Pr(t=t_{i}|h_{i})={\rm softmax}(g_{\rm time}(h_{i}))[t_{i}].$ Ultimately, the conditional probability $\Pr(v_{i}|v_{1},\dots,v_{i-1})$ given the previous observations is the product of these two distributions, representing a joint probability over POI and time: $\Pr(v_{i}|v_{1},\dots,v_{i-1})=\Pr(l=l_{i}|h_{i})\Pr(t=t_{i}|h_{i}).$
@@ 2.3.2. The objective function:
To optimize the baseline generator, given a training trajectory sequence $\mathbf{v}=(v_{1},\dots,v_{i})$, we employ cross entropy loss as our objective function. The loss is calculated as follows:
@@ 2.4. The two bottlenecks
Under the constraints of DP, it is known that the lower bound of empirical risk of the private model polynomially scales (Bassily et al., 2014, 2020) as the number of model parameters increases. In the baseline model, the size of the parameters $\theta$ expands linearly with increasing number of POIs ($n_{\rm POI}$) due to the location embedding matrix and the scoring component. Consequently, the empirical risk would polynomially escalate as $n_{\rm POI}$ increases, which leads to worse utility of the generated human mobility.
@@ 2.4.2. The number of labels
With an increase in the number of class labels (i.e., $n_{\rm POI}$), the model faces challenges in discerning more subtle feature distinctions between classes (see Section 2.3.2). This often necessitates the use of larger model architectures or extended training epochs. However, such strategies are impractical under DP-SGD constraints. In addition to the increased empirical risk due to the increased model size as mentioned above, increasing the number of training epochs results in greater privacy loss due to the composition theorem of DP (Kairouz et al., 2015). Therefore, learning difficulty caused by the large number of labels is a bottleneck of DP-SGD.
2.4.3. Discussion: impact of the discretization parameter
In many deep learning methods (Wang et al., 2023; Lian et al., 2020; Lim et al., 2022), as in this study, latitudes and longitudes are discretized using a grid. However, determining the parameters for this discretization (i.e., $w$ in this study) is not straightforward. This is because finer discretization makes the data more sparse and increases the number of parameters required for the network (location encoding). As discussed above, DP-SGD leads to a decrease in accuracy in these cases. However, finer grids have the advantage of capturing more detailed movements between grid cells, improving the quality of the training data and potentially increasing accuracy. Thus, there is a trade-off between the accuracy improvement from higher-quality training data and the accuracy decrease due to the noise required by DP. Therefore, mitigating the two bottlenecks makes it less susceptible to the negative effects of high-resolution discretization and thereby improves accuracy.
@@ 2.5. Transposed Convolution
Consider the matrix $M$ with shape $(w,w,n_{\rm dim})$, where $n_{\rm dim}$ denotes the number of channels, and each channel comprises a $w\times w$ square matrix. Similarly, let the deconvolutional kernel be defined as a matrix of shape $(2,2,n_{\rm dim})$, containing $n_{\rm dim}$ kernels, and each of these kernels is $2\times 2$ matrix denoted by ${\rm kernel}_{k}$. The 2D quad deconvolutional network expands the matrix $M$ to the matrix $M^{\prime}$ with dimensions $(2w,2w,n_{\rm dim})$ using a deconvolutional kernel. See Figure 1 for the running example. In the second layer, the kernel slides over the input of $(2\times 2)$ matrix to perform the deconvolution operation for the four grids, expanding a $2\times 2$ matrix into a $4\times 4$ matrix.